OpenStudy (anonymous):

can some one show me how to solve this? Find the sum. (8x2 + 17x) + 6x2 = 31x2 31x5 14x2 + 17x 15x2 + 17x

OpenStudy (anonymous):

all twos are supost to be exponetS!!!

OpenStudy (anonymous):

@anikit042

OpenStudy (anonymous):

@ankit042

OpenStudy (anonymous):

quack

OpenStudy (anonymous):

? can you help?

OpenStudy (ankit042):

join the like terms 8+6 =??

OpenStudy (paki):

=(8x2 + 17x) + 6x2 =14x2+17x =x(14x-17)

OpenStudy (anonymous):

ok thats 14x but do i add the exponets of 2?

OpenStudy (paki):

then 14x2+17x is ok... dont take common...

OpenStudy (anonymous):

ok well that helped alot just did 3 more problems but one one more has me stumped...